OpenStudy (anonymous):

Which set of three numbers could represent the lengths of the sides of a right triangle? A.) 7, 24 , 25 B.) 8 , 9, 10 C.) 9 , 11 , 14 D.) 15, 18 , 21

OpenStudy (jessie13):

its B

OpenStudy (anonymous):

lets check.. a^2 + b^2 = c^2 8^2 + 9^2 = 10^2 64 + 81 = 100 145 = 100 (no way....B is not the answer) Use pythagorean theorem A^2 + B^2 = C^2 To find the answer, assign your numbers letters.....doesn't really matter which ones you use for a and b ...but make sure c is the biggest number. Both sides have to equal.

OpenStudy (jessie13):

sorry its A